OpenStudy (anonymous):

Which is the same function as f(x) = ln x/3 A.) g(x) = ln x - ln 3 B.) g(x) = ln x/ln 3 c.) g(x) = ln 3 - ln x d.) g(x) = ln x + ln 3 Please explain

OpenStudy (anonymous):

Properties of logarithms say that \(\log(a/b)=\log(a)-\log(b)\). Use this property to come up with an answer.

OpenStudy (anonymous):

\[ \ln(\color{red}x/\color{blue}3) \]\[ \ln(\color{red}a/\color{blue}b) \]

OpenStudy (anonymous):

It is called pattern matching. It is fundamental part of algebra.